OpenStudy (anonymous):

Compute the flux of the vector field, F, through the surface, S. vector F= 7xi + yj + zk and S is the part of the surface z + 4x + 2y = 12 in the first octant oriented upward

OpenStudy (anonymous):

help me please @ganeshie8 :)

OpenStudy (anonymous):

help me @perl

OpenStudy (anonymous):

@eliassaab can you help me out professor

OpenStudy (surry99):

have you sketched the plane on an x, y, z coordinate system?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

@dan815 can you help me please

OpenStudy (anonymous):

The normal vector to the plane z + 4 x + 2 y = 12 is n = {4, 2, 1} and the field is {7 x, y ,z} First compute their dot product

OpenStudy (anonymous):

The dot product {4, 2, 1} . {7 x, y ,z}= 28x+ 2y + z You have to integrate (as surface integral) 28 x + 2 y + z over the part of the plane in the first octant

OpenStudy (anonymous):

Knowing that z = 12 - 4 x -2y, then 28 x + 2 y + z= 28 x + 2 y + 12 - 4 x -2y= 12 + 24 x The length of the normal vector n \(\sqrt{21}\) Can you finish it now?

OpenStudy (anonymous):

The last step is to compute this integral \[ \int_0^3 \left(\int_0^{6-2 x} \sqrt{21} (24 x+12) \, dy\right) \, dx=324 \sqrt{21} \]

OpenStudy (anonymous):

that's the same thing I got but it's wrong

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

I think it is right, unless you did not state the problem correctly

OpenStudy (anonymous):

see :/

OpenStudy (anonymous):

The problem can be wrongly programmed

OpenStudy (anonymous):

how would I use this formula on this problem

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

the answer is just 324 don't need the length. I used the flux formula

OpenStudy (anonymous):

@eliassaab